# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 1.3. Stochastic Volterra equation and stochastic heat equation
(1.29) $Y_{t}=\eta+\frac{\beta}{\sqrt{2\pi}}\int_{0}^{t}\frac{Y_{s}}{(t-s)^{\frac{1}{2}-H}}{\mathrm{d}}W_{s},$
where we replaced the constant $c$ by $\beta/\sqrt{2\pi}$ for later use. For $H\in[0,\frac{1}{2})$, the SVE (1.29) has no classic solution, as can be seen by a calculation of regularity of the terms in the integral [${}^{+}$20]. In particular, for $H\in(0,\frac{1}{2})$, the equation falls within the so-called subcritical domain, extensively explored in previous literature. In this regime, the solution has been defined in [${}^{+}$20] using regularity structures. Parallelly, the subcritical SVE has also been studied in [Com20] via rough paths for $H\in(\frac{1}{4},\frac{1}{2})$, and in [PT21] based on paracontrolled distributions for $H\in(\frac{1}{3},\frac{1}{2})$. However, for the critical SVE where $H=0$, a rigorous interpretation of the equation (1.29) remains open. Below, we will focus on the critical SVE and attempt to address this open question.
@@
(1.30) $\partial_{t}u(x,t)=\frac{1}{2}\Delta u(x,t)+\beta u(x,t)\cdot\delta_{0}(x)\dot{W}_{t},\quad u(x,0)=\eta(x),\quad(x,t)\in{\mathbbm{R}}\times{\mathbbm{R}}^{+},$
@@
(1.31) $g_{t}(x):=\frac{1}{\sqrt{2\pi t}}e^{-\frac{x^{2}}{2t}}\quad\text{with}\quad g_{0}(0):=1,$
@@
(1.32) $u(t,x)=\int_{\mathbbm{R}}g_{t}(x-y)\eta(x){\mathrm{d}}y+\beta\int_{0}^{t}\int_{\mathbbm{R}}g_{t-s}(x-y)u(s,y)\delta_{0}(y)\dot{W}_{s}{\mathrm{d}}y{\mathrm{d}}s.$
@@
Nevertheless, with the random noise term in SHE, the solution $u$ should be a distribution, and thus the equations (1.30) and (1.32) are ill-defined due to the product of two distributions $u(t,x)$ and $\delta_{0}(x)\dot{W}_{t}$. To make sense of such critical SHE, we introduce the following mollifying procedure. Let
(1.33) $\rho^{\delta}(x):=\frac{1}{\delta}\rho\Big{(}\frac{x}{\delta}\Big{)},\quad x\in{\mathbbm{R}},$
with $\rho(\cdot)$ a smooth even probability density on ${\mathbbm{R}}$ with compact support. The mollified equation
(1.34) $\begin{split}\partial_{t}u^{\delta}(t,x)=\frac{1}{2}\partial_{xx}u^{\delta}(t,x)+\beta u^{\delta}(t,x)\rho^{\delta}(x)\dot{W}_{t},\quad u^{\delta}(0,x)=\eta(x)\end{split}$
is well-defined. Using the time-reversed noise $\dot{\widetilde{W}}_{t}:=\dot{W}_{1-t}$, the solution $(u^{\delta}(t,x))_{t\in[0,1],x\in{\mathbbm{R}}}$ admits a generalised Feynman-Kac representation (see [BC95, CSZ17]) with
(1.35) $u^{\delta}(t,x)={\mathbf{E}}^{1-t,x}\Big{[}\eta(B_{1})\exp\Big{(}\beta\int_{1-t}^{1}\rho^{\delta}(B_{s}){\mathrm{d}}\widetilde{W}_{s}-\frac{1}{2}\beta^{2}\int_{1-t}^{1}(\rho^{\delta}(B_{s}))^{2}{\mathrm{d}}s\Big{)}\Big{]},$
where ${\mathbf{E}}^{1-t,x}$ denotes the expectation for a Brownian motion $B_{\cdot}$ independent of $\widetilde{W}_{\cdot}$, starting at $x$ at time $1-t$. The right-hand side of (1.35) is a natural continuous point-to-plane analogue of the directed polymer (1.25). Thereby as $\delta\downarrow 0$, a result similar to Theorem 1.7 is expected to hold.
The precise result is stated as follows. We only treat the marginal distribution at $t=1$ (and the case for general $t$ follows from Remark 1.8). With a slight abuse of notation, denote ${\mathbf{E}}^{0,x}$ and the associated probability ${\mathbf{P}}^{0,x}$ by ${\mathbf{E}}^{x}$ and ${\mathbf{P}}^{x}$, respectively. By a change of variables $(s,y)=(\delta^{2}\widetilde{s},\delta\widetilde{y})$ and by the invariance of diffusive scaling for Brownian motion, $u^{\delta}(1,x)$ is equal in distribution to the following integral,
(1.36) $\begin{split}&{\mathbf{E}}^{\frac{x}{\delta}}\Big{[}\eta(B_{1})\exp\Big{(}\beta\int_{0}^{\delta^{-2}}\rho(B_{\widetilde{s}}){\mathrm{d}}\widetilde{W}_{\widetilde{s}}-\frac{1}{2}\beta^{2}\int_{0}^{\delta^{-2}}(\rho(B_{\widetilde{s}}))^{2}{\mathrm{d}}\widetilde{s}\Big{)}\Big{]}\\ =&\int_{y\in{\mathbbm{R}}}\eta(y){\mathbf{E}}^{\frac{x}{\delta}}\Big{[}\exp\Big{(}\beta\int_{0}^{\delta^{-2}}\!\!\!\rho(B_{\widetilde{s}}){\mathrm{d}}\widetilde{W}_{\widetilde{s}}-\frac{1}{2}\beta^{2}\int_{0}^{\delta^{-2}}\!\!\!(\rho(B_{\widetilde{s}}))^{2}{\mathrm{d}}\widetilde{s}\Big{)}\Big{|}B_{\delta^{-2}}=\frac{y}{\delta}\Big{]}{\mathbf{P}}^{\frac{x}{\delta}}\Big{(}B_{\delta^{-2}}\in\frac{{\mathrm{d}}y}{\delta}\Big{)}.\end{split}$
@@
(1.37) $\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y):={\mathbf{E}}^{\frac{x}{\delta}}\Big{[}\exp\Big{(}\beta_{\delta}\int_{0}^{\delta^{-2}}\rho(B_{s}){\mathrm{d}}W_{s}-\frac{1}{2}\beta_{\delta}^{2}\int_{0}^{\delta^{-2}}(\rho(B_{s}))^{2}{\mathrm{d}}s\Big{)}\Big{|}B_{\delta^{-2}}=\frac{y}{\delta}\Big{]}\hskip 1.0pt{\mathbf{P}}^{x}(B_{1}\in{\mathrm{d}}y){\mathrm{d}}x,$
@@
(1.38) $\beta_{\delta}^{2}:=\frac{2\pi}{\log\delta^{-2}}+\frac{\theta+o(1)}{(\log\delta^{-2})^{2}},\quad\text{for some fixed}~{}\theta\in{\mathbbm{R}}.$
@@
The random measure $\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)$, defined as in (1.37), converges in distribution to the unique limit ~$\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)$ defined in Theorem 1.6, where with the mollifier $\rho(\cdot)$ given in (1.33),
(1.39) $\vartheta=\log 2-\gamma+\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})\log\Big{[}\frac{1}{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}\Big{]}\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y+\frac{\theta}{2\pi}\hskip 1.0pt.$
@@
In the proof of Theorem 1.10, we will actually show that for $\varphi\in C_{c}({\mathbbm{R}})$ and $\psi\in C_{b}({\mathbbm{R}})$, $\iint_{{\mathbbm{R}}^{2}}\varphi(x)u_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$ converges weakly to $\iint_{{\mathbbm{R}}^{2}}\varphi(x)\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$ as $\delta\to 0$.
@@
(1.40) $\widetilde{u}_{s,t}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y):={\mathbf{E}}^{\frac{s}{\delta^{2}},\frac{x}{\delta}}\Big{[}\exp\Big{(}\beta_{\delta}\int_{s\delta^{-2}}^{t\delta^{-2}}\rho(B_{s}){\mathrm{d}}W_{s}-\frac{1}{2}\beta_{\delta}^{2}\int_{s\delta^{-2}}^{t\delta^{-2}}(\rho(B_{s}))^{2}{\mathrm{d}}s\Big{)}\Big{|}B_{t\delta^{-2}}=\frac{y}{\delta}\Big{]}\hskip 1.0pt{\mathbf{P}}^{s,x}(B_{t}\in{\mathrm{d}}y){\mathrm{d}}x.$
By Remark 1.8, we have that $(\widetilde{u}^{\delta}_{s,t})_{0\leq s<t\leq 1}$ converges in finite dimensional distributions to (~$(\widetilde{\mathscr{L}}_{s,t}^{\vartheta})_{0\leq s<t\leq 1}$. Note that by (1.35)-(1.36), the solution $u^{\delta}(t,x)$ of the mollified SHE (1.34) satisfies the relation $u^{\delta}(t,x){\mathrm{d}}x=\widetilde{u}^{\delta}_{1-t,1}({\mathrm{d}}x,\eta):=\int_{y\in{\mathbbm{R}}}\widetilde{u}^{\delta}_{1-t,1}({\mathrm{d}}x,{\mathrm{d}}y)\eta(y)$. Therefore, in view of Theorem 1.10 and Remark 1.11, if the initial random function $\eta(x)$ is bounded and continuous, then the critical singular SHE (1.30) has a natural candidate solution ${\mathcal{U}}(t,{\mathrm{d}}x)$ given by
(1.41) ${\mathcal{U}}(t,{\mathrm{d}}x)=\widetilde{\mathscr{L}}_{1-t,1}^{\vartheta}({\mathrm{d}}x,\eta):=\int_{y\in{\mathbbm{R}}}\widetilde{\mathscr{L}}_{1-t,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)\eta(y).$
We then deal with the critical singular SVE (1.29) with the help of (1.35)-(1.37). Since $\eta$ is independent of $B$ and $W$, we assume without loss of generality that $\eta\equiv 1$. Decomposing $\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)$ according to the first time in $(0,1)$ that the Brownian motion $B$ visits 00, we have that
(1.42) $\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)=\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\int_{0<t<1}\Big{(}Q(x,t)\widetilde{u}_{t,1}^{\delta}(0,{\mathrm{d}}y){\mathrm{d}}x\Big{)}{\mathrm{d}}t,$
@@
(1.43) $\int_{y\in{\mathbbm{R}}}\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)=\int_{y\in{\mathbbm{R}}}\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\int_{0<t<1}\Big{[}Q(x,t)\Big{(}\int_{y\in{\mathbbm{R}}}\widetilde{u}_{t,1}^{\delta}(0,{\mathrm{d}}y)\Big{)}{\mathrm{d}}x\Big{]}{\mathrm{d}}t.$
@@
(1.44) ~$\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,\mathbbm{1}_{\mathbbm{R}})=\int_{y\in{\mathbbm{R}}}\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\int_{0<t<1}Q(x,t){\mathcal{Y}}({\mathrm{d}}t){\mathrm{d}}x,$
@@
(1.45) ${\mathcal{Y}}({\mathrm{d}}t)=\int_{t<s<1}\int_{y\in\mathbb{R}}Q(-y,1-s){\mathrm{d}}y\mathscr{L}^{\vartheta}({\mathrm{d}}t,{\mathrm{d}}s)=\sqrt{\frac{2}{\pi}}\int_{t<s<1}s^{-1/2}\mathscr{L}^{\vartheta}({\mathrm{d}}t,{\mathrm{d}}s).$
@@
Let $\beta_{\delta}$ be specified as in (1.38) and $u^{\delta}(t,x)$ be the solution of mollified SHE (1.34). Then $u^{\delta}(t,x){\mathrm{d}}x$ converges in distribution to the random measure ${\mathcal{U}}(t,{\mathrm{d}}x)$, defined in (1.41), which can be viewed as a solution of SHE (1.30). Meanwhile, $u^{\delta}(t,0){\mathrm{d}}t$ converges in distribution to a random measure ${\mathcal{Y}}({\mathrm{d}}t)$, given in (LABEL:def:cY), which can be viewed as a solution of SVE (1.29) with critical Hurst index $H=0$.
1.4. Contributions and discussions

### v3

@@ 1.3 Stochastic Volterra equation and stochastic heat equation
$Y_{t}=\eta+\frac{\beta}{\sqrt{2\pi}}\int_{0}^{t}\frac{Y_{s}}{(t-s)^{\frac{1}{2}-H}}{\mathrm{d}}W_{s},$ (1.29)
where we replaced the constant $c$ by $\beta/\sqrt{2\pi}$ for later use. For $H\in[0,\frac{1}{2})$, the SVE (1.29) has no classic solution, as can be seen by a calculation of regularity of the terms in the integral [5]. In particular, for $H\in(0,\frac{1}{2})$, the equation falls within the so-called subcritical domain, extensively explored in previous literature. In this regime, the solution has been defined in [5] using regularity structures. Parallelly, the subcritical SVE has also been studied in [18] via rough paths for $H\in(\frac{1}{4},\frac{1}{2})$, and in [33] based on paracontrolled distributions for $H\in(\frac{1}{3},\frac{1}{2})$. However, for the critical SVE where $H=0$, a rigorous interpretation of the equation (1.29) remains open. Below, we will focus on the critical SVE and attempt to address this open question.
@@
$\partial_{t}u(x,t)=\frac{1}{2}\Delta u(x,t)+\beta u(x,t)\cdot\delta_{0}(x)\dot{W}_{t},\quad u(x,0)=\eta(x),\quad(x,t)\in{\mathbbm{R}}\times{\mathbbm{R}}^{+},$ (1.30)
@@
$g_{t}(x):=\frac{1}{\sqrt{2\pi t}}e^{-\frac{x^{2}}{2t}}\quad\text{with}\quad g_{0}(0):=1,$ (1.31)
@@
$u(t,x)=\int_{\mathbbm{R}}g_{t}(x-y)\eta(x){\mathrm{d}}y+\beta\int_{0}^{t}\int_{\mathbbm{R}}g_{t-s}(x-y)u(s,y)\delta_{0}(y)\dot{W}_{s}{\mathrm{d}}y{\mathrm{d}}s.$ (1.32)
@@
Nevertheless, with the random noise term in SHE, the solution $u$ should be a distribution, and thus the equations (1.30) and (1.32) are ill-defined due to the product of two distributions $u(t,x)$ and $\delta_{0}(x)\dot{W}_{t}$. To make sense of such critical SHE, we introduce the following mollifying procedure. For $\delta>0$, let
$\rho^{\delta}(x):=\frac{1}{\delta}\rho\Big(\frac{x}{\delta}\Big),\quad x\in{\mathbbm{R}},$ (1.33)
with $\rho\in C_{c}^{\infty}({\mathbbm{R}})$ an even probability density. The mollified equation
$\begin{split}\partial_{t}u^{\delta}(t,x)=\frac{1}{2}\partial_{xx}u^{\delta}(t,x)+\beta u^{\delta}(t,x)\rho^{\delta}(x)\dot{W}_{t},\quad u^{\delta}(0,x)=\eta(x)\end{split}$ (1.34)
is well-defined. Using the time-reversed noise $\dot{\widetilde{W}}_{t}:=\dot{W}_{1-t}$, the solution $(u^{\delta}(t,x))_{t\in[0,1],x\in{\mathbbm{R}}}$ admits a generalised Feynman-Kac representation (see [8, 12]) with
$u^{\delta}(t,x)={\mathbf{E}}^{1-t,x}\Big[\eta(B_{1})\exp\Big(\beta\int_{1-t}^{1}\rho^{\delta}(B_{s}){\mathrm{d}}\widetilde{W}_{s}-\frac{1}{2}\beta^{2}\int_{1-t}^{1}(\rho^{\delta}(B_{s}))^{2}{\mathrm{d}}s\Big)\Big],$ (1.35)
where ${\mathbf{E}}^{1-t,x}$ denotes the expectation for a Brownian motion $B_{\cdot}$ independent of $\widetilde{W}_{\cdot}$, starting at time-space (1t,x)(1-t,x). The right-hand side of (1.35) is a natural continuous point-to-plane analogue of the directed polymer (1.25). Thereby as $\delta\downarrow 0$, a result similar to Theorem 1.7 is expected to hold.
[…]
$\begin{split}&{\mathbf{E}}^{\frac{x}{\delta}}\!\Big[\eta(\delta B_{\delta^{-2}})\exp\Big(\beta\int_{0}^{\delta^{-2}}\rho(B_{\widetilde{s}}){\mathrm{d}}\widetilde{W}_{\widetilde{s}}-\frac{1}{2}\beta^{2}\int_{0}^{\delta^{-2}}(\rho(B_{\widetilde{s}}))^{2}{\mathrm{d}}\widetilde{s}\Big)\Big]=\\ \int_{y\in{\mathbbm{R}}}\!\!\eta(y)&{\mathbf{E}}^{\frac{x}{\delta}}\!\Big[\exp\Big(\beta\int_{0}^{\delta^{-2}}\!\!\!\!\rho(B_{\widetilde{s}}){\mathrm{d}}\widetilde{W}_{\widetilde{s}}-\frac{1}{2}\beta^{2}\int_{0}^{\delta^{-2}}\!\!\!\!(\rho(B_{\widetilde{s}}))^{2}{\mathrm{d}}\widetilde{s}\Big)\Big|B_{\delta^{-2}}=\frac{y}{\delta}\Big]{\mathbf{P}}^{\frac{x}{\delta}}\!\Big(B_{\delta^{-2}}\in\frac{{\mathrm{d}}y}{\delta}\Big).\end{split}$ (1.36)
@@
$\begin{split}&\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y):=\\ {\mathbf{E}}^{\frac{x}{\delta}}\Big[\exp\Big(\beta_{\delta}&\int_{0}^{\delta^{-2}}\hskip-8.0pt\rho(B_{s}){\mathrm{d}}W_{s}-\frac{1}{2}\beta_{\delta}^{2}\int_{0}^{\delta^{-2}}\hskip-8.0pt(\rho(B_{s}))^{2}{\mathrm{d}}s\Big)\Big|B_{\delta^{-2}}=\frac{y}{\delta}\Big]{\mathbf{P}}^{x}(B_{1}\in{\mathrm{d}}y){\mathrm{d}}x,\end{split}$ (1.37)
@@
$\beta_{\delta}^{2}:=\frac{2\pi}{\log\delta^{-2}}+\frac{\theta+o(1)}{(\log\delta^{-2})^{2}},\quad\text{for some fixed}~\theta\in{\mathbbm{R}}.$ (1.38)
@@
The random measure $\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)$, defined as in (1.37), converges in distribution to the unique limit $\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)$ defined in Theorem 1.6, where with the mollifier $\rho(\cdot)$ given in (1.33),
$\vartheta=\log 2-\gamma+\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})\log\Big[\frac{1}{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}\Big]\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y+\frac{\theta}{2\pi}\hskip 1.0pt.$ (1.39)
@@
In the proof of Theorem 1.10, we will actually show that for $\varphi\in C_{c}({\mathbbm{R}})$ and $\psi\in C_{b}({\mathbbm{R}})$, $\iint_{{\mathbbm{R}}^{2}}\!\varphi(x)\widetilde{u}_{0,1}^{\delta}\!({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$ converges weakly to $\iint_{{\mathbbm{R}}^{2}}\!\varphi(x)\widetilde{\mathscr{L}}_{0,1}^{\vartheta}\!({\mathrm{d}}x,{\mathrm{d}}y)\psi(y)$ as $\delta\to 0$.
@@
$\begin{split}&\widetilde{u}_{s,t}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y):=\\ {\mathbf{E}}^{\frac{s}{\delta^{2}},\frac{x}{\delta}}\Big[\exp\Big(\beta_{\delta}&\int_{s\delta^{-2}}^{t\delta^{-2}}\hskip-8.0pt\rho(B_{s}){\mathrm{d}}W_{s}-\frac{1}{2}\beta_{\delta}^{2}\int_{s\delta^{-2}}^{t\delta^{-2}}\hskip-8.0pt(\rho(B_{s}))^{2}{\mathrm{d}}s\Big)\Big|B_{t\delta^{-2}}=\frac{y}{\delta}\Big]\hskip 1.0pt{\mathbf{P}}^{s,x}(B_{t}\in{\mathrm{d}}y){\mathrm{d}}x.\end{split}$ (1.40)
By Remark 1.8, we have that $(\widetilde{u}^{\delta}_{s,t})_{0\leq s<t\leq 1}$ converges in finite dimensional distributions to $(\widetilde{\mathscr{L}}_{s,t}^{\vartheta})_{0\leq s<t\leq 1}$. Note that by (1.35)-(1.36), the solution $u^{\delta}(t,x)$ of the mollified SHE (1.34) satisfies the relation $u^{\delta}(t,x){\mathrm{d}}x=\widetilde{u}^{\delta}_{1-t,1}({\mathrm{d}}x,\eta):=\int_{y\in{\mathbbm{R}}}\widetilde{u}^{\delta}_{1-t,1}({\mathrm{d}}x,{\mathrm{d}}y)\eta(y)$. Therefore, in view of Theorem 1.10 and Remark 1.11, if the initial random function $\eta(x)$ is bounded and continuous, then the critical singular SHE (1.30) has a natural candidate solution ${\mathcal{U}}(t,{\mathrm{d}}x)$ given by
${\mathcal{U}}(t,{\mathrm{d}}x)=\widetilde{\mathscr{L}}_{1-t,1}^{\vartheta}({\mathrm{d}}x,\eta):=\int_{y\in{\mathbbm{R}}}\widetilde{\mathscr{L}}_{1-t,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)\eta(y).$ (1.41)
We then deal with the critical singular SVE (1.29) with the help of (1.35)-(1.37). Since $\eta$ is independent of $B$ and $W$, we assume without loss of generality that $\eta\equiv 1$. Decomposing $\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)$ according to the first time in $(0,1)$ that the Brownian motion $B$ visits 0, we have that
$\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)=\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\int_{0<t<1}\Big(Q(x,t)\widetilde{u}_{t,1}^{\delta}(0,{\mathrm{d}}y){\mathrm{d}}x\Big){\mathrm{d}}t,$ (1.42)
@@
$\int_{y\in{\mathbbm{R}}}\!\widetilde{u}_{0,1}^{\delta}({\mathrm{d}}x,{\mathrm{d}}y)\!=\!\int_{y\in{\mathbbm{R}}}\!\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\int_{0<t<1}\!\Big[Q(x,t)\Big(\int_{y\in{\mathbbm{R}}}\widetilde{u}_{t,1}^{\delta}(0,{\mathrm{d}}y)\Big){\mathrm{d}}x\Big]{\mathrm{d}}t.$ (1.43)
@@
$\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,\mathbbm{1}_{\mathbbm{R}})=\int_{y\in{\mathbbm{R}}}\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\int_{0<t<1}Q(x,t){\mathcal{Y}}({\mathrm{d}}t){\mathrm{d}}x,$ (1.44)
@@
${\mathcal{Y}}({\mathrm{d}}t)=\int_{t<s<1}\int_{y\in\mathbb{R}}Q(-y,1-s){\mathrm{d}}y\mathscr{L}^{\vartheta}({\mathrm{d}}t,{\mathrm{d}}s)=\sqrt{\frac{2}{\pi}}\int_{t<s<1}s^{-1/2}\mathscr{L}^{\vartheta}({\mathrm{d}}t,{\mathrm{d}}s).$ (1.45)
@@
Let $\beta_{\delta}$ be specified as in (1.38) and $u^{\delta}(t,x)$ be the solution of mollified SHE (1.34). Then $u^{\delta}(t,x){\mathrm{d}}x$ converges in distribution to the random measure ${\mathcal{U}}(t,{\mathrm{d}}x)$, defined in (1.41), which can be viewed as a solution of the SHE (1.30). Meanwhile, $u^{\delta}(t,0){\mathrm{d}}t$ converges in distribution to a random measure ${\mathcal{Y}}({\mathrm{d}}t)$, given in (1.45), which can be viewed as a solution of the SVE (1.29) with critical Hurst index $H=0$.
1.4 Contributions and discussions